OpenStudy (anonymous):

solve 1/4x+2= 5 for x

OpenStudy (anonymous):

1/(4x+2)=5 =>(1/5)=4x+2 =>.2=4x+2 =>4x=.2-2 =>4x=(-1.8) =>x=-(.45)

OpenStudy (anonymous):

it cant be 45

OpenStudy (anonymous):

Its -(0.45)

OpenStudy (johnweldon1993):

Is the question like \[\frac{ 1 }{ 4 }x + 2 = 5\] or \[\frac{ 1 }{ 4x + 2 } = 5\]

OpenStudy (anonymous):

the 1st one john

OpenStudy (anonymous):

Oh i've answered for 2nd one

OpenStudy (anonymous):

lol oh

OpenStudy (johnweldon1993):

Okay...then you want to isolate 1/4x first...so you subtract 2 from both sides 1/4x = 3 Now...to cancel out 1/4x ...you multiply both sides by 4 1/4 times 4 = 1...so it would cancel it out So we have 1/4x * 4 = 3 * 4 x = 12

OpenStudy (anonymous):

first one is quiet easy Jhon Already answered that :)

OpenStudy (anonymous):

Thanks john can i get help in another one ?

OpenStudy (johnweldon1993):

absolutely :)

OpenStudy (anonymous):

Solve for H: A = L • W • H

OpenStudy (johnweldon1993):

So this says A = LWH And we need to solve for H (just rearrange this equation to solve for H....H = ...) So remember how we "isolated" 1/4x above....we got rid of everything else that was around it? we would do the same here.... A = LWH To isolate H....we need to divide both sides by LW LW / LW = 1...so that will cancel it out on that right side...and all that will be left is H... A = LWH -------- LW LW A/LW = H or written correctly \[H = \frac{ A }{ LW }\]

OpenStudy (johnweldon1993):

Lol hope you got to see that and understand it before Openstudy went out :)